solve the following question

  1. dy/dx+3y=3x2e-x
1
Expert's answer
2021-07-19T10:34:06-0400

Let us solve the following differential equation "\frac{dy}{dx}+3y=3x^2e^{-x}." For this let us multply both parts by "e^{3x}:" "e^{3x}\frac{dy}{dx}+3e^{3x}y=3x^2e^{-x}e^{3x}." It follows that "(e^{3x}y)'=3x^2e^{2x}." Then "e^{3x}y=3\int x^2e^{2x}dx=|u=x^2,dv=e^{2x}dx,du=2xdx, v=\frac{1}{2}e^{2x}|=\frac{3}{2}x^2e^{2x}-3\int xe^{2x}dx=|u=x,dv=e^{2x},du=dx,v=\frac{1}{2}e^{2x}|=\frac{3}{2}x^2e^{2x}-\frac{3}{2}xe^{2x}+\frac{3}{2}\int e^{2x}dx=\frac{3}{2}x^2e^{2x}-\frac{3}{2}xe^{2x}+\frac{3}{4}e^{2x}+C."

It follows that the general solution is "y=e^{-3x}(\frac{3}{2}x^2e^{2x}-\frac{3}{2}xe^{2x}+\frac{3}{4}e^{2x}+C)=\frac{3}{2}x^2e^{-x}-\frac{3}{2}xe^{-x}+\frac{3}{4}e^{-x}+Ce^{-3x} ."